1/1

Christian Morgenstern,

The Werewolf

tradukita de Alexander Gross

A Werewolf, troubled by his name, Left wife and brood one night and came To a hidden graveyard to enlist The aid of a long-dead philologist.

“Oh sage, wake up, please don’t berate me,” He howled sadly, “Just conjugate me.” The seer arose a bit unsteady Yawned twice, wheezed once, and then was ready.

“Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense, And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

“I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied. The scholar paused–again he tried:

“A ‘Will-be-wolf?’ It’s just too long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly wrong! Such words are wounds beyond all suture– I’m sorry, but you have no future.”

The Werewolf knew better–his sons still slept At home, and homewards now he crept, Happy, humble, without apology For such folly of philology.

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

http://www.poezio.net/version?poem-id=825&version-id=,1657